1.00 mL of a 2.90x10-4 M solution of oleic acid (C18H34O2) is diluted with 9.00 mL of petroleum ether (C6H14), forming solution A. Then 2.00 mL of solution A is diluted with 8.00 mL of petroleum ether, forming solution B. What is the concentration of solution B?


1
Expert's answer
2021-05-19T03:06:33-0400

We will use dilution formula two times

For solution A

M1V1=M2V2

 2.90x10-4 *1=9M2

M2= 0.32x10-4 

FOr solution B

 0.32x10-4 *2=8*M2

M2=  0.08x10-4